Grade 11Physical Chemistry

The molality of a solution obtained by mixing 750 ml of 0.5 0.5 molarity of HCL with 250 ml of 2 molarity of HCL will be

Rohit kr patel
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Final molarity would be =>
M1V1 = M2V2
Therefore,
0.5*750 + 250*2 = M(1000)
M = (375 + 500)/1000  => 875/1000 => 0.875 M
This would be the final molarity
And to calculate the molality U need to provide the density of solution.